ganeshie8 (ganeshie8):

[Solved by @dg2 ] find the last digit of this sum \(\LARGE \sum \limits_{n=1}^{100} n^n \)

OpenStudy (jack1):

1

ganeshie8 (ganeshie8):

nice try haha! but no : http://www.wolframalpha.com/input/?i=%5Csum+%5Climits_%7Bn%3D1%7D%5E100+n%5En

OpenStudy (jack1):

;)

OpenStudy (ikram002p):

\(1^1+2^2+3^3+4^4+....+ 100^{100} \mod 10 \)

ganeshie8 (ganeshie8):

yep !

ganeshie8 (ganeshie8):

this might be hard as i was not bale to find any good solution anywhere online...

OpenStudy (ikram002p):

brb

OpenStudy (anonymous):

20

OpenStudy (anonymous):

so 0

ganeshie8 (ganeshie8):

how did u get 20 ?

OpenStudy (anonymous):

1^1-last digit1 2^2-4 3^3=7 4^4=6 5^5=5 6^6=6 7^7=3 8^8=6 9^9=9 10^10=0

ganeshie8 (ganeshie8):

interesting, could you explain a bit more... :)

OpenStudy (anonymous):

thinking!.give me some more minutes

ganeshie8 (ganeshie8):

i think you're on right track, take your time.. .

OpenStudy (anonymous):

power cycle of 1^1 is 1 ok

ganeshie8 (ganeshie8):

yes

OpenStudy (anonymous):

power cycle of 3 is 3,9,7,1 power cycleof 9 is 7,9,3,1 .......................7 is same as above

OpenStudy (anonymous):

you know power cycle of 2 is ,6,8,10,12

OpenStudy (anonymous):

4 is 6 always

ganeshie8 (ganeshie8):

it should be : 2,4,8,6,2... right?

OpenStudy (ikram002p):

cool ^^

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

5^5 is always 5

ganeshie8 (ganeshie8):

oh yes!

OpenStudy (anonymous):

6^6,and 8 s4*2 cycles 10 is 0

OpenStudy (anonymous):

so add all we get the answer i think

ganeshie8 (ganeshie8):

brilliant ! i think we can simply add first 10 powers and multiply them by 10

OpenStudy (anonymous):

1 is 10 times=10 2 is 10 times 2,4,6,8etc =80

ganeshie8 (ganeshie8):

1^1-last digit1 2^2-4 3^3=7 4^4=6 5^5=5 6^6=6 7^7=3 8^8=6 9^9=9 10^10=0 ---------------- 7

ganeshie8 (ganeshie8):

7*10 = 70 so units place is 0 fantastic work @dg2 !!

OpenStudy (ikram002p):

well iwould like to do sum mod 2 shold be = sum mod 5 sum mod 2 1+0+1+0+1+0+1+0 .....( 50 times) mod 2 so 50 mod 2 = 0 mos 2 = 0 mod 10 mmm for sum mod 5 1+2+3+4+5 +1+2+3+4+5 + ....... ( without ordering )** mod 5 = 5 k mod 5 =0 ** trick with fermat mmm

OpenStudy (anonymous):

10+80+60+60+50+60+60+50+60+0=490 like this also i think

OpenStudy (ikram002p):

wait i need to work more on this part for sum mod 5 1+2+3+4+5 +1+2+3+4+5 + ....... ( without ordering )** mod 5 = 5 k mod 5 =0 ** trick with fermat mmm

OpenStudy (anonymous):

oh i dont understand that mod what @ikram002p is using.and idk about mod .give some examples

OpenStudy (ikram002p):

:3

ganeshie8 (ganeshie8):

@ikram002p are you using chinese remaidner thm ?

OpenStudy (ikram002p):

not really , will i used this identity first if a,b primes x=m mod a x=m mod b then x= m mod ab 10=2*5 its more simple to work with primes instead of composite numbers

ganeshie8 (ganeshie8):

I see... you want to work the remainders by first dividing the sum by 2 and 5 ?

OpenStudy (ikram002p):

however , \(\LARGE \sum \limits_{n=1}^{100} n^n \mod 2 \) since terms are odd/even then it would be like this 1+0+1+0+1+0+.... +0 mod 2=50 mod 2 = 0 mod 2 exactly ganesh

ganeshie8 (ganeshie8):

that way mod2 is easy, but working mod5 is very painful ^^

OpenStudy (ikram002p):

now lets see fermats apply that \(a^4=1 \mod 5\)

OpenStudy (ikram002p):

^^

ganeshie8 (ganeshie8):

Wow! so that kills 25 terms in the sum, what aboutthe remaining 75 terms ?

ganeshie8 (ganeshie8):

Fermat reduces all exponents of form 4k to 1, right ?

OpenStudy (ikram002p):

from 1 to 3 we will check by hand when power n is :- n=4k ,then mod =0 25 elements the others will be from 1 to 0 xD the mod sum will be 1+2+3+4+0+1+2 ( without ordering ) mod 5 :)

OpenStudy (ikram002p):

@dg2 method is also awesome :3

OpenStudy (anonymous):

but i dont understand

OpenStudy (ikram002p):

from 1 to 4 **

OpenStudy (ikram002p):

wait made a typo there i meant mod when its 4k is 1 when its 5k is zero

OpenStudy (anonymous):

help me.some what undertood when divided by 10 we have two factors 2*5 but when divided by 2 we get 50 number as remainder zero,5 means 20 number as zero we got 60 numbers as remainder zero

OpenStudy (anonymous):

a^4=1 mod 5?

OpenStudy (ikram002p):

this is a theorm , Fermat little theorms ok ill write down in neat way :P which im not good in :P

OpenStudy (anonymous):

okay

OpenStudy (anonymous):

i think u should also hav patience

OpenStudy (ikram002p):

hehe i also dont have that as well

OpenStudy (anonymous):

create it

ganeshie8 (ganeshie8):

lol okay they spend one full semister studying number theory @dg2

OpenStudy (ikram002p):

:3 patience can be neither created nor destroyed, but can change form

ganeshie8 (ganeshie8):

haha! the key thing to notice is this : when you divide by 2, you get only two numbers as remainders : {0, 1}

OpenStudy (anonymous):

yes

ganeshie8 (ganeshie8):

so performing division by 2 is a piece of cake : 1^1 + 2^2 + 3^3 + 4^4 + 5^5 + .... + 100^100 mod 2 is same as : 1 + 0 + 1 + 0 + .... + 0 mod 2

OpenStudy (ikram002p):

who they ? sry Os laging to me

ganeshie8 (ganeshie8):

which is same as 50 mod 2

ganeshie8 (ganeshie8):

which is same as 0 mod 2

OpenStudy (anonymous):

then 60 ones and 40ones

OpenStudy (ikram002p):

:o

ganeshie8 (ganeshie8):

Is that clear why the reaminder of the entre sum is 0 when divided by 2 ?

OpenStudy (anonymous):

srry 60 zeros,40 ones

OpenStudy (ikram002p):

50 zeroes ,50 ones

ganeshie8 (ganeshie8):

make it 50 zeroes, 50 ones :P

OpenStudy (anonymous):

when divided by 5 we get 20 zeros

ganeshie8 (ganeshie8):

every odd number leaves a remainder 1 every even number leaves a remainder 0 50 odd numbers, 50 even numbers between 1 & 100 so 50 1's and 50 0's

ganeshie8 (ganeshie8):

thats right !

ganeshie8 (ganeshie8):

we will be left with 80 terms in the sum, which @ikram002p has worked by using Fermat and other advanced number theory tricks which i am also not able to understand fully

OpenStudy (anonymous):

:-):-):-):)

OpenStudy (ikram002p):

ohh ,ok ill go through it step by step :) tbh when u get through it its the same as dg did so we wanna find :- \(\LARGE \sum \limits_{n=1}^{100} n^n \mod 5\)

OpenStudy (anonymous):

he told he will explain neatly

OpenStudy (ikram002p):

Fermat set that for any number a >2 , then \(a^4 =1 \mod 5\)

OpenStudy (ikram002p):

so \(a^{4k}=1 \mod 5\) thus \(a^{4k+1}=a \mod 5\) \(a^{4k+2}=a^2 \mod 5\) \(a^{4k+3}=a^3 \mod 5\)

OpenStudy (ikram002p):

ok so far ?

OpenStudy (anonymous):

okay

OpenStudy (anonymous):

\[a^{p-1} \equiv 1 \pmod{p}.\]

OpenStudy (ikram002p):

PS , we do not count a =5n sence mod would be zero @ganeshie8 reaso these step for me i understand it but cant explain hehe like we would have a it self might be 4k 4k+1 4k+2 4k+3

OpenStudy (anonymous):

formula right?

OpenStudy (ikram002p):

yeah Fermat :)

OpenStudy (anonymous):

its quite confusing .i will see it later.after working some problems using fermat

OpenStudy (ikram002p):

ok so the mod sum well be like this :-( this might clear doubts ) \(\ \sum \limits_{n=1}^{100} n^n=1+2^2+3^3+1 +0 + 6^2+7^3+1+9+0 +11^3 +1+13+14^2 +15^3+1+17+18^2+19^3+0+....+ ect xD\)

OpenStudy (ikram002p):

\(\ \sum \limits_{n=1}^{100} n^n=1+2^2+3^3+1 +0 + 6^2+7^3\) \(+1+9+0 +11^3 +1+13+14^2 +15^3+1+17+18^2+19^3+0+....+ \) ect xD

OpenStudy (ikram002p):

mmm thats it -.- we would have 1+2+3+4+0+.....(without ordering ) =10 k mod 5 = 0

ganeshie8 (ganeshie8):

nice xD

OpenStudy (ikram002p):

<3

OpenStudy (anonymous):

good